(c) Determine the interval and radius of convergence for the following power series.
-(4x - 12-11 11 0 (5 marks)
Answer #1

Radius of convergence =

\frac{1}{24}

Interval=\left(\frac{5}{24},\frac{7}{24}\right)

Solution- E 6 (40-1)^- n n=6 on = 6 (40)n-1 n (40) ant = Ст+ ) nti Ch+) (Ч-1), 17 9 lim nyo On an htl 6h (404)nt n 1241 - 611 || २५ च - । - । । २५ ५ २५ 5 २५ 1 / 1 २५ Intuival of Convergence 5 २५ २५

If any doubt kindly comment. Thank you.